OpenStudy (anonymous):

Let S be the subspace of x+y+z=3 in R^3. Let V be (1,0,1). Find distance of v to s.

OpenStudy (anonymous):

can you find the Normal to x+y+z=3 ?

OpenStudy (anonymous):

sorry typo is x+Y+Z=0

OpenStudy (anonymous):

I not sure about the subspace basis of x+y+z=0 is it (0,0,0) or (1,0,-1) or other?

OpenStudy (anonymous):

the normal to the plane is (1,1,1) right ?

OpenStudy (anonymous):

o so the distance should be 1?